### v2

Abstract
Wide-angle videos in few-shot action recognition (FSAR) effectively express actions within specific scenarios. However, without a global understanding of both subjects and background, recognizing actions in such samples remains challenging because of the background distractions. Receptance Weighted Key Value (RWKV), which learns interaction between various dimensions, shows promise for global modeling. While directly applying RWKV to wide-angle FSAR may fail to highlight subjects due to excessive background information. Additionally, temporal relation degraded by frames with similar backgrounds is difficult to reconstruct, further impacting performance. Therefore, we design the CompOund SegmenTation and Temporal REconstructing RWKV (Otter). Specifically, the Compound Segmentation Module (CSM) is devised to segment and emphasize key patches in each frame, effectively highlighting subjects against background information. The Temporal Reconstruction Module (TRM) is incorporated into the temporal-enhanced prototype construction to enable bidirectional scanning, allowing better reconstruct temporal relation. Furthermore, a regular prototype is combined with the temporal-enhanced prototype to simultaneously enhance subject emphasis and temporal modeling, improving wide-angle FSAR performance. Extensive experiments on benchmarks such as SSv2, Kinetics, UCF101, and HMDB51 demonstrate that Otter achieves state-of-the-art performance. Extra evaluation on the VideoBadminton dataset further validates the superiority of Otter in wide-angle FSAR.
Codehttps://github.com/wenbohuang1002/Otter
1 Introduction
The difficulties of video collection and labeling complicates traditional data-driven training based on fully labeled datasets. Fortunately, few-shot action recognition (FSAR) improves learning efficiency and reduces the labeling dependency by classifying unseen actions from extremely few video samples. Therefore, FSAR has diverse real-world applications, including health monitoring and motion analysis (yan2023feature; wang2023openoccupancy). However, recognizing similar actions under regular viewpoint is a non-trival problem in FSAR. For instance, distinguishing “indoor climbing” and “construction working” is challenging, as subjects exhibit similar actions against a wall. To mitigate this issue, wide-angle videos provide contextual background, such as a “climbing wall” or a “construction site”, expressing actions within specific scenarios more accurately. According to established definitions (lai2021correcting; zhang2025madcow), wide-angle videos with a greater field of view (FoV) are widespread [^1] This work adopts the widely accepted definition of wide-angle as FoV exceeding 80.. FoV estimation (lee2021ctrl; hold2023perceptual) on popular FSAR benchmarks further reveals that approximately 35% of samples per dataset fall into this category, yet remain unexplored.
Figure 1: Smaller subject proportion (red circles) and degraded temporal relation (red dotted lines) both contribute to background distractions in wide-angle FSAR. As a result, wide-angle samples are more challenging to recognize compared with regular samples.
On the other hand, effectively modeling wide-angle videos remains a critical issue due to the difficulty of accurately interpreting both subjects and background content. Recent success in recurrent model-based architectures has led to methods such as Receptance Weighted Key Value (RWKV) (peng2023rwkv; peng2024eagle), which demonstrate strong performance in global modeling across various tasks by enabling token interaction through linear interpolation, thereby expanding the receptive field and efficiently capturing subject–background dependencies.
To seamlessly apply RWKV in wide-angle FSAR, two key challenges remain, primarily due to background distractions, as illustrated in Figure 1. Challenge 1: Lack of primary subject highlighting in RWKV. As shown in the “snowboarding” examples, the primary subject occupies a smaller proportion in wide-angle frames. When RWKV is directly applied for global feature extraction, it tends to capture massive secondary background information “snow” rather than the primary subject “athlete”. Since the background serves as contextual information while the subject is crucial for determining feature representation, this reversal of primary and secondary information may lead to potential misclassification. Challenge 2: Absence of temporal relation reconstruction in RWKV. Temporal relation plays a significant role in FSAR, primarily in perceiving action direction and aligning frames. From the “snowboarding” example, we observe that abundant background information in similar frames obscures the evolution of primary subject “athlete”, causing temporal relation degraded in wide-angle samples. However, RWKV focuses on global modeling but lacks the capability to reconstruct temporal relation, increasing the difficulty of recognizing wide-angle samples.
Although current attempts achieve promising results (fu2020depth; wang2023molo; perrett2021temporal; huang2024soap; wang2022hybrid; xing2023revisiting), few works address the two aforementioned challenges simultaneously. Therefore, we propose the CompOund SegmenTation and Temporal REconstructing RWKV (Otter), which highlights subjects and restores temporal relations in wide-angle FSAR. To be specific, we devise the Compound Segmentation Module (CSM) to adaptively segment each frame into patches and highlight the subject before feature extraction. This enables RWKV to focus on the subject rather than being overwhelmed by secondary background information. We further design the Temporal Reconstruction Module (TRM), integrated into temporal-enhanced prototype construction to perform bidirectional feature scanning across frames, enabling RWKV to reconstruct temporal relations degraded in wide-angle videos. Additionally, we combine a regular prototype with a temporal-enhanced prototype to simultaneously achieve subject highlighting and temporal relation reconstruction. This strategy significantly improves the performance of wide-angle FSAR.
To the best of our knowledge, the proposed Otter is the first attempt of utilizing RWKV for wide-angle FSAR. The core contribution is threefold.
  • The CSM is introduced to highlight the primary subject in RWKV. It segments each frame into multiple patches, learns adaptive weights from each patch to highlight the subject, and then reassembles the patches in their original positions. This process enables more effective detection of inconspicuous subjects in wide-angle FSAR.
  • The TRM is designed to reconstruct temporal relations in RWKV. It performs bidirectional scanning of frame features and reconstructs the temporal relation via a weighted average of the scanning results for the temporal-enhanced prototype. This module mitigates temporal relation degradation in wide-angle FSAR.
  • The state-of-the-art (SOTA) performance achieved by Otter is validated through extensive experiments on prominent FSAR benchmarks, including SSv2, Kinetics, UCF101, and HMDB51. Additional analyses on wide-angle VideoBadminton dataset emphasize superiority of Otter, particularly in wide-angle FSAR.
2 Related works
2.1 Few-Shot Learning
Few-shot learning, which aims to classify unseen classes using extremely limited samples, is a crucial area in the deep learning community (fei2006one). It encompasses three main paradigms: augmentation, optimization, and metric-based. Augmentation-based methods (hariharan2017low; wang2018low; zhang2018metagan; chen2019image; li2020adversarial) address data scarcity by generating synthetic samples to augment the training set. In contrast, optimization-based methods (finn2017model; ravi2017optimization; rusu2018meta; jamal2019task; rajeswaran2019meta) focus on modifying the optimization process to enable efficient fine-tuning with few samples. Among these approaches, the metric-based paradigm (snell2017prototypical; oreshkin2018tadam; sung2018learning; hao2019collect; wang2020cooperative) is the most widely adopted in practical applications due to its simplicity and effectiveness. Specifically, these methods construct class prototypes and perform classification by the similarity between query features and class prototypes using learnable metrics.
[…]
2.2 Few-Shot Action Recognition
Metric-based meta-learning is the mainstream paradigm in FSAR due to its simplicity and effectiveness. This approach embeds support features into class prototypes to represent various classes. Most methods rely on temporal alignment to match queries with prototypes. For example, the dynamic time warping (DTW) algorithm is used in OTAM for similarity calculation (cao2020few). Subsequent works, including ITANet (zhang2021learning), T2AN (li2022ta2n), and STRM (thatipelli2022spatio), further optimize temporal alignment. To focus more on local features, TRX (perrett2021temporal), HyRSM (wang2022hybrid), SloshNet (xing2023boosting), SA-CT (zhang2023importance), and Manta (huang2025manta) employ fine-grained or multi-scale modeling. Additionally, models are enhanced with supplementary information such as depth (fu2020depth), optical flow (wanyan2023active), and motion cues (wang2023molo; wu2022motion; huang2024soap). Despite achieving satisfactory performance, They are unable to address challenges in wide-angle FSAR simultaneously.
2.3 RWKV Model
The RWKV model is initially proposed for natural language processing (NLP) (peng2023rwkv; peng2024eagle), combining the parallel processing capabilities of Transformers with the linear complexity of RNNs. This fusion enables RWKV to achieve efficient global modeling with reduced memory usage and accelerated inference speed following data-driven training. Building on this foundation, the vision-RWKV (VRWKV) model is developed for computer vision tasks and has demonstrated notable success (duan2024vision). Additionally, numerous studies have explored integrating RWKV with Diffusion or CLIP, achieving remarkable results in various domains (fei2024diffusion; gu2024rwkv; he2024pointrwkv; yuan2024mamba). However, the potential of RWKV in wide-angle FSAR remains unexplored.
3 Methodology
3.1 Problem Definition
Following settings in previous literature (cao2020few; perrett2021temporal), three parts including training set $\mathcal{D}_{\text{train}}$, validation set $\mathcal{D}_{\text{val}}$, and testing set $\mathcal{D}_{\text{test}}$ without overlap ($\mathcal{D}_{\text{train}}\cap\mathcal{D}_{\text{val}}\cap\mathcal{D}_{\text{test}}=\varnothing$) are divided from datasets. Each part is further split into two non-overlapping sets including support $\mathcal{S}$ with at least one labeled sample of each class and query $\mathcal{Q}$ with all unlabeled samples ($\mathcal{S}\cap\mathcal{Q}=\varnothing$). The aim of FSAR is to classify samples from $\mathcal{Q}$ into one class of $\mathcal{S}$. A large number of few-shot tasks are randomly selected and combined from $\mathcal{D}_{\text{train}}$. We define few-shot setting as $N$-way $K$-shot from $\mathcal{S}$ with $N$ classes, $K$ samples in each class.
Successive $F$ frames are uniformly extracted from a video each time. The $k^{\text{th}}$ ($k=1,\cdots,K$) sample of the $n^{\text{th}}$ ($n=1,\cdots,N$) class of $\mathcal{S}$ is defined as $S^{n,k}$ and randomly selected sample from $\mathcal{Q}$ is denoted as $Q^{r}$ ($r\in\mathbb{Z}^{+}$).
$S^{n,k}$ $=\left[s_{1}^{n,k},\dots,s_{F}^{n,k}\right]\in\mathbb{R}^{F\times C\times H\times W},$ (1)
$Q^{\gamma}$ $=\left[q_{1}^{\gamma},\dots,q_{F}^{\gamma}\right]\in\mathbb{R}^{F\times C\times H\times W},$
in which $F$, $C$, $H$, and $W$ represent frames, channels, height, and width, respectively.
3.2 Overall Architecture
We demonstrate the overall architecture of Otter via a simple 3-way 3-shot example in Figure 2. The following two main components of Otter are built from specific combinations of core units (§ 3.3). At the first stage of motion segmentation, CSM works for highlighting subjects before feature extracting via backbone (§ 3.4). TRM is introduced in the second stage of prototype 1 (temporal-enhanced) construction, reconstructing the temporal relation (§ 3.5). Prototype 2 (regular) construction is the third stage, retaining subject emphasis (§ 3.5). Finally, distances calculated from weighted average of two prototypes are employed in cross-entropy loss $\mathcal{L}_{\text{ce}}$. In order to further distinguish class prototypes, the prototype similarities serve as $\mathcal{L}_{\text{p}}^{1}$ and $\mathcal{L}_{\text{p}}^{2}$. The weighted combination of three loss including $\mathcal{L}_{\text{ce}}$, $\mathcal{L}_{\text{p}}^{1}$ and $\mathcal{L}_{\text{p}}^{2}$ is the training objective $\mathcal{L}_{\text{total}}$ (§ 3.6).
3.3 Core Units
(a) Spatial Mixing
(b) Time Mixing
(c) Channel Mixing
[…]
To be specific, Spatial Mixing (Figure 3(a)) is designed to aggregate features from different spatial locations. Let $r_{t}$, $k_{t}$, $v_{t}$, and $g_{t}$ denote the $t^{\text{th}}$ features of $R$, $K^{*}$, $V$, and $G$, respectively. This design allows the model to capture dependencies across different regions of the image, thereby enhancing its ability to model global spatial features.
$\vartriangle_{t}$ $=W_{\vartriangle}\cdot\mathrm{Q}\text{-}\mathrm{Shift}_{\vartriangle}\left(x\right)$ (2)
$=W_{\vartriangle}\cdot\left[x+\left(1-\mu_{\vartriangle}\right)\odot x^{\prime}\right],\forall{\vartriangle}\in\left\{r,k^{*},v,g\right\},$
$x^{\prime}_{\left[h^{\prime},w^{\prime}\right]}$ $=\mathrm{Concat}\left(x_{\left[h^{\prime}-1,w^{\prime},0:C/4\right]},x_{\left[h^{\prime}+1,w^{\prime},C/4:C/2\right]},\right$.
$\quad\left.x_{\left[h^{\prime},w^{\prime}-1,C/2:3C/4\right]},x_{\left[h^{\prime},w^{\prime}+1,3C/4:C\right]}\right),$
where $\mu$ is a learnable vector for the calculation of $R$, $K^{*}$, and $V$ while $\mathrm{Concat}\left(\cdot\right)$ means concatenate operation. “:” separates the start and end index. Row and column index of $x$ are denoted by $h^{\prime}$ and $w^{\prime}$. Then attention result $\left(wk^{*}v\right)_{t}$ is calculated according to the following definition.
$\left(wk^{*}v\right)_{t}$ $=\mathrm{Bi}\text{-}\mathrm{WK^{*}V}\left(K^{*},V\right)_{t}$ (3)
$=\frac{\sum\nolimits_{i=0,i\neq t}^{t-1}{e^{-\left(\left|t-i\right|-1\right)\cdot w+k^{*}_{i}}v_{i}+e^{u+k^{*}_{t}}v_{t}}}{\sum\nolimits_{i=0,i\neq t}^{t-1}{e^{-\left(\left|t-i\right|-1\right)\cdot w+k^{*}_{i}}+e^{u+k^{*}_{t}}}},$
$W$ is determined by vector $w$. After combining with $r_{t}$ and $g_{t}$, the $o^{\text{th}}$ feature of output $O$ can be calculated as
$o_{t}=\sigma\left(g_{t}\right)\odot\mathrm{Norm}\left(r_{t}\otimes\left(wk^{*}v\right)_{t}\right),$ (4)
in which $\sigma\left(\cdot\right)$ denotes activation function while $\mathrm{Norm}\left(\cdot\right)$ represents normalization.
As illustrated Figure 3(b), we observe that the main discrepancies between Time Mixing and Spatial Mixing are $\vartriangle_{t}$ and $\mathrm{WK^{*}V}\left(\cdot\right)$. The former one can be defined as
$\vartriangle_{t}=W_{\vartriangle}\cdot\left[x_{t}+\left(1-\mu_{\vartriangle}\right)\odot x_{t}-1\right],\forall{\vartriangle}\in\left\{r,k^{*},v,g\right\},$ (5)
while the latter can be written as
$\left(wk^{*}v\right)_{t}$ $=\mathrm{WK^{*}V}\left(K^{*},V\right)_{t}$ (6)
$=\frac{\sum\nolimits_{i=0,i\neq t}^{t-1}{e^{-\left(t-i-1\right)\cdot w+k^{*}_{i}}v_{i}+e^{u+k^{*}_{t}}v_{t}}}{\sum\nolimits_{i=0,i\neq t}^{t-1}{e^{-\left(t-i-1\right)\cdot w+k^{*}_{i}}+e^{u+k^{*}_{t}}}},$
After achieving $O$ with the same way, the combination of current and past states enable long-term modeling.
In order to capture dependencies between multiple dimensions of input, Channel Mixing (Figure 3(c)) mixes information from various channels by $R$ and $V$, as
$O=\sigma_{r}\left(R\right)\odot\sigma_{v}\left(V\right).$ (7)
$\sigma_{r}\left(\cdot\right)$ and $\sigma_{v}\left(\cdot\right)$ means two difference kinds of activation function applied for $R$ and $V$.
3.4 Motion Segmentation
Compound Segmentation Module (CSM)
Figure 4: The structure of Compound Segmentation Module (CSM).
As demonstrated in Figure 4, each frame is segmented into $HW/p^{2}$ patches with $\mathrm{Seg}\left(\cdot,\cdot\right)$. Using random frames $s,q\in\mathbb{R}^{C\times H\times W}$ from $S^{n,k},Q^{r}$ as simple examples.
$\vartriangle^{p}=\mathrm{Seg}\left(\vartriangle,p\right)\in\mathbb{R}^{C\times p\times p},\forall{\vartriangle}\in\left\{s,q\right\}.$ (8)
$H$ and $W$ must be divisible by $p$. The operations of Spatial Mixing, Time Mixing, and Channel Mixing can be written as $\mathrm{S}\text{-}\mathrm{Mix}\left(\cdot\right)$, $\mathrm{T}\text{-}\mathrm{Mix}\left(\cdot\right)$, and $\mathrm{C}\text{-}\mathrm{Mix}\left(\cdot\right)$, respectively. The output $\vartriangle^{\alpha}$ of $\mathrm{S}\text{-}\mathrm{Mix}\left(\cdot\right)$ is connected with the input $\vartriangle^{p}$ for capturing region associations of patches, as
$\vartriangle^{\alpha}=\left[\mathrm{S}\text{-}\mathrm{Mix}\left(\vartriangle^{p}\right)\oplus\vartriangle^{p}\right]\in\mathbb{R}^{C\times p\times p},\forall{\vartriangle}\in\left\{s,q\right\}.$ (9)
The activation function $\sigma\left(\cdot\right)$ in $\mathrm{S}\text{-}\mathrm{Mix}\left(\cdot\right)$ is $\mathrm{Sigmoid}\left(\cdot\right)$. Through the same method of connection with $\vartriangle^{\alpha}$, the output $\vartriangle^{\beta}$ of $\mathrm{T}\text{-}\mathrm{Mix}\left(\cdot\right)$ can be achieved.
$\vartriangle^{\beta}=\left[\mathrm{C}\text{-}\mathrm{Mix}\left(\vartriangle^{\alpha}\right)\oplus\vartriangle^{\alpha}\right]\in\mathbb{R}^{C\times p\times p},\forall{\vartriangle}\in\left\{s,q\right\},$ (10)
where the $\sigma_{r}\left(\cdot\right)$ and $\sigma_{v}\left(\cdot\right)$ of $\mathrm{C}\text{-}\mathrm{Mix}\left(\cdot\right)$ are $\mathrm{Sigmoid}\left(\cdot\right)$ and $\mathrm{Relu}\left(\cdot\right)$. Following C3-STISR (zhao2022c3), learnable weights $lw^{\vartriangle}\in\mathbb{R}^{C\times p\times p}$ can be achieved from $\vartriangle^{p}$ and $\vartriangle^{\beta}$ via convolution $\mathrm{Conv}\left(\cdot\right)$ and residual connection.
$lw^{\vartriangle}=\mathrm{Sigmoid}\left[\mathrm{Conv}\left(\vartriangle^{\beta}\right)\oplus\vartriangle^{p}\right],\forall{\vartriangle}\in\left\{s,q\right\}.$ (11)
[…]
$\dot{\vartriangle}=\mathrm{RT}\left(\dots,lw^{\vartriangle}\odot\vartriangle^{\beta},\dots\right)\in\mathbb{R}^{C\times H\times W},\forall{\vartriangle}\in\left\{s,q\right\}.$ (12)
According to (9) and (10), the final outputs $\hat{\vartriangle}$ ($\forall{\vartriangle}\in\left\{s,q\right\}$) of CSM are calculated via $\mathrm{S}\text{-}\mathrm{Mix}\left(\cdot\right)$ and $\mathrm{C}\text{-}\mathrm{Mix}\left(\cdot\right)$. We place each $\hat{\vartriangle}$ in its raw position for residual connection with inputs $\hat{S}^{n,k},\hat{Q}^{\gamma}$, thereby achieving subject highlighting.
Feature Extraction
$D$-dimensional features $S^{n,k}_{f},Q^{\gamma}_{f}\in\mathbb{R}^{F\times D}$ are extracted by sending $\hat{S}^{n,k},\hat{Q}^{\gamma}$ into backbone $f_{\theta}\left(\cdot\right):\mathbb{R}^{C\times H\times W}\mapsto\mathbb{R}^{D}$.
3.5 Prototype Construction
Temporal Reconstruction Module (TRM)
[…]
$\tilde{\vartriangle}=\left[\vartriangle+\mathrm{Avg}\left(\grave{\vartriangle},\acute{\vartriangle}\right)\right]\in\mathbb{R}^{F\times D},\forall{\vartriangle}\in\left\{S^{n,k}_{f},Q^{\gamma}_{f}\right\}.$ (14)
After the TRM, temporal relation is recovered.
Prototype and Distance
$P^{n}_{1}$ is prototype of the $n^{\text{th}}$ support class, being achieved via average calculation of $\tilde{S}_{f}^{n,k}$:
$P_{1}^{n}={\frac{1}{K}}\sum_{k=1}^{K}{\tilde{S}_{f}^{n,k}}\in\mathbb{R}^{F\times D}.$ (15)
The distance between $\tilde{Q}^{\gamma}_{f}$ and $P_{1}^{n}$ is $D_{1}$.
$D_{1}=\left\|P_{1}^{n}-\tilde{Q}_{f}^{\gamma}\right\|.$ (16)
For further distinguishing classes of the prototype $P_{1}$, we apply the sum of cosine similarity function $\mathrm{Sim}\left(\cdot,\cdot\right)$ for $\mathcal{L}^{1}_{\text{P}}$:
$\mathcal{L}_{\mathrm{P}}^{1}=\sum_{n\neq n^{\prime}}{\mathrm{Sim}\left(P_{1}^{n},P_{1}^{n^{\prime}}\right)},\left(P_{1}^{n},P_{1}^{n^{\prime}}\right)\in P_{1}.$ (17)
The prototype 2 is constructed without TRM. Therefore, the $n^{\text{th}}$ support prototype $P^{n}_{2}$ can be computed from $S^{n,k}_{f}$. Then the corresponding distance $D_{2}$ between $Q^{\gamma}_{f}$ and $P_{2}^{n}$ can also be achieved. After the same cosine similarity calculation, $\mathcal{L}^{2}_{\text{P}}$ is applied for differentiating classes of $P_{2}$.
3.6 Training Objective
The distance $D$ between $n^{\text{th}}$ class and $Q^{\gamma}_{f}$ is the weighted mean value of $D_{1}$ and $D_{2}$ with weight $\omega$. Therefore, the predicted label $\tilde{y}_{Q}^{j}\in\tilde{Y}_{Q}$ of query is
$\tilde{y}_{Q}^{j}=\underset{n}{\mathrm{argmin}}\left(D\right),D=\sum_{i=1}^{2}{\omega_{i}D_{i}}.$ (18)
$\tilde{y}_{Q}^{j}$ and the ground truth $y_{Q}^{j}\in Y_{Q}$ are applied in cross-entropy loss $\mathcal{L}_{\text{ce}}$ calculation.
$\mathcal{L}_{\text{ce}}=-\frac{1}{N}\sum_{j=1}^{N}{y_{Q}^{j}\log\left(\tilde{y}_{Q}^{j}\right)}.$ (19)
The training objective $\mathcal{L}_{\text{total}}$ is the combination of $\mathcal{L}_{\text{ce}}$, $\mathcal{L}^{1}_{\text{P}}$, and $\mathcal{L}^{2}_{\text{P}}$ under weight factor $\lambda$ as:
$\mathcal{L}_{\text{total}}=\lambda_{0}\mathcal{L}_{\text{ce}}+\lambda_{1}\mathcal{L}^{1}_{\text{P}}+\lambda_{2}\mathcal{L}^{2}_{\text{P}},$ (20)
4 Experiments
4.1 Experimental Configuration
  Methods Reference Pre-Backbone SSv2 Kinetics UCF101 HMDB51
1-shot 5-shot 1-shot 5-shot 1-shot 5-shot 1-shot 5-shot
STRM (thatipelli2022spatio) CVPR’22 ImageNet-RN50 N/A 68.1 N/A 86.7 N/A 96.9 N/A 76.3
SloshNet (xing2023revisiting) AAAI’23 ImageNet-RN50 46.5 68.3 N/A 87.0 N/A 97.1 N/A 77.5
SA-CT (zhang2023importance) MM’23 ImageNet-RN50 48.9 69.1 71.9 87.1 85.4 96.3 61.2 76.9
GCSM (yu2023multi) MM’23 ImageNet-RN50 N/A N/A 74.2 88.2 86.5 97.1 61.3 79.3
GgHM (xing2023boosting) ICCV’23 ImageNet-RN50 54.5 69.2 74.9 87.4 85.2 96.3 61.2 76.9
[1pt/1pt] STRM (thatipelli2022spatio) CVPR’22 ImageNet-ViT N/A 70.2 N/A 91.2 N/A 98.1 N/A 81.3
SA-CT (zhang2023importance) MM’23 ImageNet-ViT N/A 66.3 N/A 91.2 N/A 98.0 N/A 81.6
TRX (perrett2021temporal) CVPR’21 ImageNet-RN50 53.8 68.8 74.9 85.9 85.7 96.3 83.5 85.5
HyRSM (wang2022hybrid) CVPR’22 ImageNet-RN50 54.1 68.7 73.5 86.2 83.6 94.6 80.2 86.1
MoLo (wang2023molo) CVPR’23 ImageNet-RN50 56.6 70.7 74.2 85.7 86.2 95.4 87.3 86.3
SOAP (huang2024soap) MM’24 ImageNet-RN50 61.9 85.8 86.1 93.8 94.1 99.3 86.4 88.4
Manta (huang2025manta) AAAI’25 ImageNet-RN50 63.4 87.4 87.4 94.2 95.9 99.2 86.8 88.6
[1pt/1pt] MoLo (wang2023molo) CVPR’23 ImageNet-ViT 61.1 71.7 78.9 95.8 88.4 97.6 81.3 84.4
SOAP (huang2024soap) MM’24 ImageNet-ViT 66.7 87.2 89.9 95.5 96.8 99.5 89.3 89.8
Manta (huang2025manta) AAAI’25 ImageNet-ViT 66.2 89.3 88.2 96.3 97.2 99.5 88.9 88.8
[1pt/1pt] MoLo (wang2023molo) CVPR’23 ImageNet-ViR 60.9 71.8 79.1 95.7 88.2 97.5 81.2 84.6
SOAP (huang2024soap) MM’24 ImageNet-ViR 66.4 87.1 89.8 95.8 96.6 99.1 88.8 89.7
Manta (huang2025manta) AAAI’25 ImageNet-ViR 66.5 89.2 88.1 96.1 96.7 99.2 88.7 89.5
AmeFu-Net (fu2020depth) MM’20 ImageNet-RN50 N/A N/A 74.1 86.8 85.1 95.5 60.2 75.5
MTFAN (wu2022motion) CVPR’22 ImageNet-RN50 45.7 60.4 74.6 87.4 84.8 95.1 59.0 74.6
AMFAR (wanyan2023active) CVPR’23 ImageNet-RN50 61.7 79.5 80.1 92.6 91.2 99.0 73.9 87.8
Lite-MKD (liu2023lite) MM’23 ImageNet-RN50 55.7 69.9 75.0 87.5 85.3 96.8 66.9 74.7
[1pt/1pt] Lite-MKD (liu2023lite) MM’23 ImageNet-ViT 59.1 73.6 78.8 90.6 89.6 98.4 71.1 77.4
[1pt/1pt] Lite-MKD (liu2023lite) MM’23 ImageNet-ViR 59.1 73.7 78.5 90.5 89.7 97.9 71.2 77.5
Otter Ours ImageNet-RN50 64.7 88.5 90.5 96.4 96.8 99.2 88.1 89.8
[1pt/1pt] Otter Ours ImageNet-ViT 67.2 89.9 91.8 97.3 97.7 99.4 89.9 90.6
[1pt/1pt] Otter Ours ImageNet-ViR 67.1 89.8 91.7 96.8 97.5 99.3 89.5 90.5
Table 1: Comparison ($\uparrow$ Acc. %) on ResNet-50 (ImageNet-RN50), ViT-B (ImageNet-ViT), and VRWKV-B (ImageNet-ViR) are separated by dashed line. Bold texts denotes the global best results while Underline texts serve as the local best. From top to bottom, the whole table is divided into three parts including RGB-based, multimodal, and our Otter. In the first two parts, $\star$” represents our implementation with the same setting. “N/A” indicates not available.
Data Processing
Temporal-related SSv2 (goyal2017something), spatial-related Kinetics (carreira2017quo), UCF101 (kay2017kinetics), and HMDB51 (kuehne2011hmdb) are most frequently-used benchmark datasets for FSAR. A wide-angle dataset VideoBadminton (li2024benchmarking) is employed for evaluating real-world performance. In order to prove the effectiveness of our Otter, the sampling intervals setting of decoding videos are each 1 frame. Based on widely-used data split (zhu2018compound; cao2020few; zhang2020few), $\mathcal{D}_{\text{train}}$, $\mathcal{D}_{\text{val}}$, and $\mathcal{D}_{\text{test}}$ ($\mathcal{D}_{\text{train}}\cap\mathcal{D}_{\text{val}}\cap\mathcal{D}_{\text{test}}=\varnothing$) are divided from each dataset. Then further split of support $\mathcal{S}$ and query $\mathcal{Q}$ are executed for FSAR.
According to TSN (wang2016temporal), each frame are sized into $3\times 256\times 256$ while $F$ of successive frames is set to 8. $3\times 224\times 224$ random crops and horizontal flipping data augmentation is applied during training while only the center crop is utilized in testing. As an exception, horizontal flipping is absent in SSv2 because of many actions with horizontal direction such as “Pulling S from left to right222“S” means “something”.”.
Implementation Details and Evaluation Metrics
Standard 5-way 1-shot and 5-shot setting are adopted for FSAR. We select ResNet-50, ViT-B, VMamba-B, and VRWKV-B with ImageNet pre-trained weights initialization as our backbone. The dimension $D$ of features is 2048.
The larger SSv2 are trained with 75,000 tasks while other datasets only require 10,000 tasks. SGD optimization for training is applied with initial learning rate $10^{-3}$. The $\mathcal{D}_{\text{val}}$ determines hyper-parameters such as distance weight ($\omega_{1}=\omega_{2}=0.5$), weight factor of loss $\lambda$ ($\lambda_{0}=0.8,\lambda_{1}=\lambda_{2}=0.1$) and patch size ($p=56$). Average accuracy of 10,000 random tasks from $\mathcal{D}_{\text{test}}$ is recorded during testing stage. Experiments are most conducted on a server with two 32GB NVIDIA Tesla V100 PCIe GPUs.
4.2 Comparison with Various Methods
We implement many methods under the same setting for fair comparison with Otter. The average accuracy ($\uparrow$ higher indicates better) is illustrated in Table 1.
ResNet-50 Methods
Using SSv2 under 1-shot setting as representative results, we find that Otter outperforms the current SOTA method Manta which focuses on long sub-sequences from 63.4% to 64.7%. A similar improvement can also be discovered in other datasets with different shots.
ViT-B Methods
The larger model capacity makes ViT-B perform better than ResNet-50. We observe that the previous SOTA performance is achieved by SOAP or Manta. Being similar with ResNet-50, Otter reveals superior performance, surpassing previous methods.
VRWKV-B Methods
As an emerging model, VRWKV-B can efficiently extract feature form promising regions association. Compared with other backbones, we observe that the overall trend in performance has no significant changes. The proposed Otter focus on improving wide-angle samples, achieving new SOTA performance.
4.3 Essential Components and Factors
Key Components
In order to analyze the effect of key components in Otter, we conduct experiments with only CSM, TRM, and both of them. As demonstrated in Table 2, we observe that CSM and TRM both improve the performance. In our design, CSM highlights subject within wide-angle frames before feature extraction. Then TRM reconstructs the degraded temporal relations. Two modules operate successively and complement each other, indicating that full Otter achieves optimal performance.
  CSM TRM SSv2 Kinetics
1-shot 5-shot 1-shot 5-shot
54.6 69.2 78.1 85.3
61.3 85.6 89.4 94.8
59.5 83.4 87.8 92.7
64.7 88.5 90.5 96.4
Table 2: Comparison ($\uparrow$ Acc. %) of key components.
Patch Design in CSM
A deeper research on patch design in CSM is indicated in Table 3. It is obvious that the performance is increasing with more fine grained design segmentation (less $p$). If $p$ is further reduced to 28, the performance will have a decline. We also consider multi-scale patch configurations and observe that p=56p{=}56 consistently performs better. This may be attributed to the fact that multi-scale design introduces redundant features. Therefore, we adopt p=56p{=}56 with $4{\times}4$ segmentation in our patch design.
     $p$ SSv2       Kinetics
1-shot 5-shot 1-shot 5-shot
$p=224$ 62.7 86.4 87.7 94.6
$p=112$ 63.6 87.1 89.5 95.2
$p=56$ 64.7 88.5 90.5 96.4
$p=28$ 64.1 87.9 90.2 95.8
$p\in\left\{28,56\right\}$ 64.2 88.1 90.1 96.1
$p\in\left\{56,112\right\}$ 63.7 87.9 89.6 95.8
Table 3: Comparison ($\uparrow$ Acc. %) of patch design in CSM.
Direction Design in TRM
[…]
1-shot 5-shot 1-shot 5-shot
63.2 87.3 89.7 95.7
60.6 85.2 89.1 94.2
64.7 88.5 90.5 96.4
Table 4: Comparison ($\uparrow$ Acc. %) of direction design in TRM.
4.4 Wide-Angle Evaluation
Performance on Wide-Angle Dataset
In order to evaluate Otter on wide-angle scenario, we employ VideoBadminton dataset with all wide-sample samples for testing. Form the results in Table 5, it is Otter that obviously far ahead of other methods without specific design for wide-angle samples. Owing to highlighted subject and reconstructed temporal relation, Otter mitigates background distractions. Therefore, the performance on challenging wide-angle samples is significantly improved.
[…]
1-shot 5-shot 1-shot 5-shot
MoLo 60.2 64.5 58.9 61.7
SOAP 63.5 66.9 60.1 63.1
Manta 64.1 67.1 62.1 65.3
Otter 71.2 75.8 69.5 72.6
[…]
CAM Visualization
In Figure 6, subjects are inconspicuous and similar background makes temporal relation degradation. From the CAM results without Otter, the focuses of model are mostly in the background while the subject in distance is entirely ignored. When being equipped with Otter, most of focuses are transferred to subjects and background is not completely overlooked. Compared with only focusing on the subject nearby, Otter can capture both of subjects playing badminton. These prove that Otter helps the model better understand “smash”, an action that requires interaction between two subjects, mitigating background distractions and achieving better performance in wide-angle FSAR.
Figure 6: CAM of “smash” from VideoBadminton. O: Original, w/: with Otter, w/o: without Otter.
Various FoV
To rigorously evaluate Otter on wide-angle samples, frames with varying FoV are essential. Given that FoV is primarily determined by complementary metal oxide semiconductor (CMOS) size and lens focal length (liao2023deep), we utilized PQDiff (zhang2024continuous) for outpainting magnification ($U_{\text{m}}$) and introduced the distortion factor ($U_{\text{d}}$) in the VideoBadminton dataset to simulate diverse CMOS sizes and focal lengths. This approach results in five distinct FoV levels, with higher levels indicating a wider FoV. As indicated in Figure 7, we observe that recent methods all have a drastic downward trend with the increasing level of FoV. Although our Otter is also negatively influenced, the downward trend is much more stable, revealing the outstanding performance of wide-angle FSAR.
(a) 1-shot
(b) 5-shot
Figure 7: Comparison ($\uparrow$ Acc. %) with various FoV levels.
5 Conclusion
In this work, we propose Otter which is specially designed against background distractions of wide-angle FSAR. Otter highlights subjects in each frames via adaptive segmentation and enhancement of CSM. Temporal relation degradation caused by too many frames with similar background is reconstructed by bidirectional scanning of TRM. Otter achieves new SOTA performance on several widely-used datasets. Further studies demonstrate the competitiveness of our proposed method, especially for mitigating background distractions of wide-angle FSAR. We hope this work will inspire upcoming research in FSAR community.
Supplementary Materials
In the supplementary material, we provide:
  • Extra Study of Key Components (mentioned in § 4.3)
  • Additional Wide-Angle Evaluation (mentioned in § 4.4)
  • Robustness Study
  • Computational Complexity
Appendix A Extra Study of Key Components
A.1 Study on RWKV-4 and RWKV-5/6
Currently, RWKV-4 (peng2023rwkv) and RWKV-5/6 (peng2024eagle) are released official versions. The main discrepancy is additional gate $G$ mechanism in RWKV-5/6 for the control of information flow. In order to compare the performance, we conduct experiments with three key components under various basis. The results are demonstrated in Table I. We find that applying RWKV-5/6 performs better than components based on RWKV-4. Therefore, we select the updated RWKV-5/6 as the basis of our proposed Otter.
     S-Mix T-Mix C-Mix SSv2 Kinetics
1-shot 5-shot 1-shot 5-shot
R-4 R-4 R-4 64.0 87.5 89.2 94.3
R-5/6 R-4 R-4 64.2 87.4 89.5 94.5
R-4 R-5/6 R-4 64.1 87.6 89.1 94.7
R-4 R-4 R-5/6 64.0 87.4 89.4 94.4
R-5/6 R-5/6 R-4 64.2 87.8 90.0 96.1
R-5/6 R-4 R-5/6 64.4 88.1 90.1 95.7
R-4 R-5/6 R-5/6 64.2 87.9 89.7 95.5
R-5/6 R-5/6 R-5/6 64.7 88.5 90.5 96.4
Table I: Comparison ($\uparrow$ Acc. %) between RWKV-4 (R-4) and RWKV-5/6 (R-5/6).
A.2 Study on Learnable Weights
[…]
1-shot 5-shot 1-shot 5-shot
61.8 85.2 85.7 91.1
63.8 87.9 89.7 95.8
62.1 86.6 89.4 95.1
64.7 88.5 90.5 96.4
Table II: Comparison ($\uparrow$ Acc. %) of learnable weights.
Loss Design
In the loss design, we fix $\mathcal{L}_{\text{ce}}$ as the primary loss for classification and reveal experiments in Table III. As auxiliary loss, both $\mathcal{L}^{1}_{\text{P}}$ and $\mathcal{L}^{1}_{\text{P}}$ combined with $\mathcal{L}_{\text{ce}}$ can improve the performance via further distinguishing similar classes of prototype. The simultaneous use of the three losses can obtain the best performance of wide-angle FSAR. Therefore, $\mathcal{L}_{\text{ce}}$, $\mathcal{L}^{1}_{\text{P}}$, and $\mathcal{L}^{1}_{\text{P}}$ are necessary in Otter.
    $\mathcal{L}_{\text{ce}}$ $\mathcal{L}^{1}_{\text{P}}$ $\mathcal{L}^{2}_{\text{P}}$       SSv2       Kinetics
1-shot 5-shot 1-shot 5-shot
63.3 84.8 89.8 95.5
63.4 88.0 90.1 95.7
64.7 88.5 90.5 96.4
Table III: Comparison ($\uparrow$ Acc. %) of loss design.
A.3 Study on Loss Weight Factors
The training objective is the combination of $\mathcal{L}_{\text{ce}}$, $\mathcal{L}^{1}_{\text{P}}$, and $\mathcal{L}^{2}_{\text{P}}$ with loss weight factors $\lambda$. Experiments are conducted and the results are illustrated in Table IV. As a role primarily used for classification, $\lambda_{0}$ for $\mathcal{L}_{\text{ce}}$ should not be less than 0.5. Considering the similar function of $\mathcal{L}^{1}_{\text{P}}$ and $\mathcal{L}^{2}_{\text{P}}$, $\lambda_{1}$ and $\lambda_{2}$ should be equal. The performance is improved with the increasing $\lambda_{0}$ but begins to decline when $\lambda_{0}>0.8$. The above results confirm the loss weight factors.
     $\lambda_{0}$ $\lambda_{1}$ $\lambda_{2}$       SSv2       Kinetics
1-shot 5-shot 1-shot 5-shot
0.50 0.25 0.25 62.9 87.6 89.6 95.6
0.60 0.20 0.20 64.1 88.0 89.9 95.9
0.70 0.15 0.15 64.3 88.2 90.3 96.2
0.80 0.10 0.10 64.7 88.5 90.5 96.4
0.90 0.05 0.05 64.4 88.4 90.2 96.2
Table IV: Comparison ($\uparrow$ Acc. %) of loss factors.
A.4 Study on Various Types of Prototype
There are three types of prototype construction including attention-based calculation $\mathrm{Attn}\left(\cdot\right)$ (wang2022hybrid), query-specific prototype $\mathrm{Q}\text{-}\mathrm{Sp}\left(\cdot\right)$ (perrett2021temporal), and averaging calculation $\mathrm{Avg}\left(\cdot\right)$ (huang2025manta). Experiments about compatibility of Otter and prototype construction is conducted in Table V. Although $\mathrm{Attn}\left(\cdot\right)$ and $\mathrm{Q}\text{-}\mathrm{Sp}\left(\cdot\right)$ with extra calculation achieve advanced performance in their work, the fitness with our Otter is not the best. Therefore, we select simple $\mathrm{Avg}\left(\cdot\right)$ as our prototype.
  Prototype SSv2 Kinetics
1-shot 5-shot 1-shot 5-shot
$\mathrm{Attn}\left(\cdot\right)$ 63.9 87.1 89.3 94.9
$\mathrm{Q}\text{-}\mathrm{Sp}\left(\cdot\right)$ 64.5 88.5 90.1 96.0
$\mathrm{Avg}\left(\cdot\right)$ 64.7 88.5 90.5 96.4
Table V: Comparison ($\uparrow$ Acc. %) of various prototype types.
Appendix B Additional Wide-Angle Evaluation
B.1 Details of Wider FoV Simulation
From previous definition (liao2023deep), FoV is only determined by camera CMOS size ($H_{\text{c}}\times W_{\text{c}}$) and lens focal length ($L_{\text{f}}$). Related calculation is written as
$FoV=2\mathrm{arctan}\left(\frac{\vartriangle}{2L_{\text{f}}}\right),\forall\vartriangle\in\left(H_{\text{c}},W_{\text{c}}\right).$ (I)
Image size is positively correlated with the CMOS size, while the distortion is negatively correlated with the focal length (hu2022miniature). Therefore, directly applying larger outpainting magnification ($U_{\text{m}}$) and introducing larger distortion factor ($U_{\text{d}}$) can simulate wider FoV. A group of simulation with five various levels is provided in Figure I. We observe that a wider FoV means more background. Meanwhile, distortion is more exaggerated. Wide-angle datasets always correct them for stable training. Re-adding distortion makes wide-angle FSAR more challenging.
(a) Lv.0
(b) Lv.1
(c) Lv.2
(d) Lv.3
(e) Lv.4
Figure I: Examples with various wide FoV levels. To be specific, each level is a combination of $U_{\text{m}}$ and $U_{\text{d}}$.
B.2 Temporal Relation
According to OTAM (cao2020few), DTW scores calculated from two sequences ($\downarrow$ lower indicates better) can reflect the quality of temporal relation via alignment degree. The curves are shown in Figure II. We observe that models equipped with Otter converge much faster than those without Otter under any few-shot setting. The convergence points for the 5-shot are much earlier due to the increased number of training samples. Under the 1-shot setting of FSAR, the DTW curves without Otter even do not converge under Lv.1 or 2 FoV, indicating a more time-consuming training. Therefore, it is evident that Otter can effectively reconstruct temporal relations of wide-angle FSAR.
(a) 1-shot
(b) 5-shot
Figure II: $\downarrow$ DTW scores under 1-shot setting during training on VideoBadminton. In specific, “Lv.1, w/” denotes model equipped with Otter and samples with Lv.1 FoV.
B.3 T-SNE Visualization
From the t-SNE (van2008visualizing) revealed in Figure III, the wide-angle actions are hard to be separated and clustered well without any assistance. Samples with Lv.4 FoV simulation are scattered everywhere. The above observation prove the difficulties in wide-angle FSAR. On the contrary, Otter clusters samples from same class and scatters others better. Although these special samples with 100% expanding magnification are located at the edge of each class, the cluster condition of them is much better.
(a) w/o Otter
(b) w/ Otter
Figure III: T-SNE visualization of five action classes in support from Kinetics (25-shot). Blue: “ice skating”, Orange: “snowboarding”, Green: “paragliding”, Red: “skateboarding”, Purple: “crossing river”. Dots with black borders are samples with Lv.4 FoV simulation.
B.4 Additional CAM Visualization
Additional CAM visualization for wide-angle samples are provided in Figure IV. Taking “crossing river” as an example, it is evident that the model without Otter focuses on “forests” due to their larger proportion in the frames. Although subject “Jeep” is included, recognition is inevitably interfered with by the background. In contrast, Otter accurately highlights the subject while not completely ignoring the background, thereby achieving better performance. This focus pattern is consistent across the other two examples. These CAM visualization demonstrate that Otter mitigates background distractions, helping models better understand challenging actions in wide-angle scenario.
(a) crossing river
(b) paragliding
(c) driving tractor
Figure IV: Additional CAM of “crossing river”, “paragliding”, and “driving tractor” from Kinetics dataset.
Appendix C Robustness Study
In order to explore robustness of Otter, we select two groups of noise added into $\mathcal{D}_{\text{test}}$ of FSAR. The first group is task-based including sample-level and frame-level noise for simulating unexpected circumstances during data collection. As revealed in Figure V, another group is visual noise such as zoom, Gaussian, rainy, and light noise, for simulating different shooting situations. Specifically, zoom frames are imposed by variation in optimal zoom while Gaussian noise is related to digital issues of hardware. Changeable weather and lighting conditions result in rainy and light noise.
(a) O
(b) Z
(c) G
(d) R
(e) L
Figure V: Different kinds of noise. In specific, O, Z, G, R, and L denote original frames, zoom, Gaussian, rainy, and light noise, respectively.
C.1 Sample-Level Noise
Wide-angle samples from other classes may be mixed into a particular class. Correcting sample-level noise is time-consuming and laborious. Therefore, directly testing wide-angle FSAR on sample-level noise can reflect the robustness of a method. The experimental results are indicated in Table VI. It is obvious that the introduce of sample-level has negative impacts on the performance of wide-angle FSAR. The results decline with the increasing ratio of sample-level noise. However, we find that the robustness of our proposed Otter is better than other recent methods.
  Datasets Methods Sample-Level Noise Ratio
0% 10% 20% 30% 40%
SSv2 MoLo 72.5 70.5 68.2 66.4 64.1
SOAP 87.3 85.1 83.0 80.8 78.7
Manta 89.6 87.6 86.2 83.1 80.9
Otter 90.2 89.4 88.2 86.6 85.5
Kinetics MoLo 87.5 85.1 83.4 80.8 78.1
SOAP 95.9 94.2 92.1 89.7 87.5
Manta 96.1 94.2 91.9 90.1 87.8
Otter 98.4 97.5 96.2 95.0 93.8
Table VI: Comparison ($\uparrow$ Acc. %) with sample-level noise under 5-way 10-shot setting.
C.2 Frame-Level Noise
Multiple irrelevant frames mixed into wide-angle samples are called as frame-level noise. Serving as a unexpected situation of data collection, robustness of methods can also be reflected by frame-level noise. From the results in Table VII, we observe that the performance of wide-angle FSAR is harmed with the increasing number of noisy frames. The reason for this phenomenon is that frame-level noise further disorganizes subjects and temporal relation. Under the circumstance, our Otter still reveals stable performance, reflecting better robustness of frame-level noise.
  Datasets Methods Noisy Frame Numbers
0 1 2 3 4
SSv2 MoLo 72.5 69.3 66.5 63.3 59.6
SOAP 87.3 84.1 80.9 78.0 75.6
Manta 89.6 86.4 83.2 80.4 77.3
Otter 90.2 89.0 88.2 87.2 86.0
Kinetics MoLo 87.5 84.3 81.5 78.0 75.3
SOAP 95.9 93.1 90.2 87.6 84.1
Manta 96.1 93.0 89.7 86.8 83.7
Otter 98.4 97.1 95.7 94.5 93.2
Table VII: Comparison ($\uparrow$ Acc. %) with frame-level noise under 5-way 10-shot setting.
C.3 Visual-Based Noise
Visual-based noise challenges the robustness of a method. Therefore, we add each type of visual-based noise to 25% samples for creating more complex wide-angle FSAR tasks. As shown in Table VIII, the zoom noise has the largest negative impact on the performance. Other types of visual-based noise more or less harm the results. However, we observe that our Otter can keep the SOTA performance under those challenging environment. These phenomena in wide-angle FSAR reflect the better robustness of the proposed Otter.
  Datasets Methods Visual-Based Noise Type
O Z G R L
SSv2 MoLo 72.5 70.0 70.3 69.7 69.8
SOAP 87.3 84.7 84.0 84.6 86.1
Manta 89.6 87.5 88.7 88.8 87.4
Otter 90.2 89.6 89.6 89.3 89.0
Kinetics MoLo 87.5 85.2 86.3 86.7 85.9
SOAP 95.9 93.6 94 94.4 93.9
Manta 96.1 93.9 95.0 95.1 94.8
Otter 98.4 97.9 98.0 97.7 97.8
Table VIII: Comparison ($\uparrow$ Acc. %) with various types of 25% visual-based noise under 5-way 10-shot setting.
C.4 Cross Dataset Testing
In real-world scenario, various data distributions are exist. Therefore, we applying the cross dataset method (training and testing on various datasets) for the simulation of different data distributions. SSv2 and Kinetics with three no-overlapping set are utilized. Then overlapping classes of $\mathcal{D}_{\text{train}}$ and $\mathcal{D}_{\text{test}}$ from different datasets are further removed. From the results revealed in Table IX, despite cross-dataset setting degrades the performance, Otter can keep ahead of other methods. This trend similar with the regular test setting highlights the robustness of Otter.
[…]
1-shot 5-shot 1-shot 5-shot
MoLo 53.7 (56.6) 68.7 (70.7) 71.5 (74.2) 83.2 (85.7)
SOAP 60.0 (61.9) 84.5 (85.8) 84.1 (86.1) 91.1 (93.8)
Manta 61.5 (63.4) 86.4 (87.4) 86.3 (87.4) 91.8 (94.2)
Otter 63.1 (64.7) 86.7 (88.5) 89.2 (90.5) 94.0 (96.4)
[…]
C.5 Any-Shot Testing
In real-world application, ensuring shot number of each class equal is challenging. In order to create a more authentic testing environment for robustness, we apply the any-shot setup ($1\leqslant K\leqslant 5$). From the results demonstrated in Table X, we observe that the performance of Otter defeats other methods, reflecting our Otter has a better robustness for applications in real-world scenario.
  Methods SSv2 Kinetics
MoLo 64.6 $\pm\text{1.5}$ 80.2 $\pm\text{1.8}$
SOAP 73.8 $\pm\text{1.4}$ 89.1 $\pm\text{1.3}$
Manta 75.2 $\pm\text{1.1}$ 90.6 $\pm\text{1.3}$
Otter 77.4 $\pm\text{0.6}$ 93.6 $\pm\text{0.7}$
Table X: Comparison ($\uparrow$ Acc. %) with 95% confidence interval of 5-way any-shot setting.
Appendix D Computational Complexity
D.1 Inference Speed
To evaluate the model under practical conditions with limited resources, we conducted 10,000 tasks using a single 24GB NVIDIA GeForce RTX 3090 GPU on a server. From results demonstrated in Table XI, we find that the inference speed of MoLo and SOAP is slow because of Transformer with high computational complexity. On the contrary, Mamba-based Manta and RWKV-based Otter is much faster than previous Transformer-based methods. Considering the accuracy of classification, the proposed Otter is more suitable for practical applications.
  Methods SSv2 Kinetics
1-shot 5-shot 1-shot 5-shot
MoLo 7.83 8.02 7.64 8.14
SOAP 7.44 7.86 7.21 7.72
Manta 4.25 4.61 4.42 4.56
Otter 4.13 4.24 4.35 4.48
Table XI: Inference speed ($\downarrow$ hour) with 10,000 random tasks on single 24GB NVIDIA GeForce RTX 3090 GPU.
D.2 Major Tensor Changes
[…]
  Operation Input Input Size Output Output Size
[…]
$\mathrm{RT}$ $lw^{\vartriangle}\odot\vartriangle^{\beta}$ $pp] ˙\dot{\vartriangle} [FF, CC,$ $H$, $W$]
CSM $S^{c,k},Q^{\gamma}$ [$F$, $C$, $H$, $W$] $\hat{S}^{c,k},\hat{Q}^{\gamma}$ [$F$, $C$, $H$, $W$]
$f_{\theta}$ $\hat{S}^{c,k},\hat{Q}^{\gamma}$ [$F$, $C$, $H$, $W$] $S^{c,k}_{f},Q^{\gamma}_{f}$ [$F$, $D$]
[…]
D.3 Pseudo-Code
The primary computational burden lies in the Compound Segmentation Module (CSM). For complexity analysis, the related pseudo-code is listed in Algorithm 1. Considering the low computational complexity of core units including $\mathrm{S}\text{-}\mathrm{Mix}\left(\cdot\right)$, $\mathrm{T}\text{-}\mathrm{Mix}\left(\cdot\right)$, and $\mathrm{C}\text{-}\mathrm{Mix}\left(\cdot\right)$ in RWKV, the functions $\mathrm{Seg}(\cdot,\cdot)$ and $\mathrm{RT}(\dots,\cdot,\dots)$ form the main structure with nested loops. Both the inner and outer loops have a computational complexity of $O(p)$. Consequently, the total complexity of the CSM is $O(p^{2})$. Given the determined size and single-scale design of $p$ (p=56)(p={56}), the additional computational burden introduced by Otter is negligible, ensuring its usability in real-world applications.
Input: $\vartriangle\in\mathbb{R}^{F\times C\times H\times W},\forall{\vartriangle}\in\left\{S^{c,k},Q^{\gamma}\right\}$, $NP~\left(H\mid NP,W\mid NP,NP\in\mathbb{Z}^{+}\right)$.
Output: $\hat{\vartriangle}\in\mathbb{R}^{F\times C\times H\times W},\forall{\vartriangle}\in\left\{S^{c,k},Q^{\gamma}\right\}$.
1 $\hat{\vartriangle}_{1}\leftarrow\varnothing$;
2 if $H\%NP==0$ & $W\%NP==0$ then
3 $Block\leftarrow\left(H/NP,W/NP\right)$ ;
4 if $Block\left[0\right]==Block\left[1\right]$ then
5    $p_{H}\leftarrow Block\left[0\right]$ ;
6    $p_{W}\leftarrow Block\left[1\right]$ ;
7    for each $i\in\left[0,p_{H}\right]$ do
8       for each $j\in\left[0,p_{W}\right]$ do
          /* $\mathrm{Seg}\left(\cdot,\cdot\right)$ Start */
9          $s\leftarrow\left[i\times p_{H},j\times p_{W}\right]$;
10          $e\leftarrow\left[\left(i+1\right)\times p_{H},\left(j+1\right)\times p_{W}\right]$;
11          $p\leftarrow\vartriangle\left[:,:,s\left(0\right):e\left(0\right),s\left(1\right):e\left(1\right)\right]$;
12          $p1\leftarrow\textnormal{{S-Mix}}\left(p\right)\oplus p$ ;
13          $p2\leftarrow\textnormal{{C-Mix}}\left(p1\right)\oplus p1$ ;
14          $lw\leftarrow\sigma\left[\textnormal{{Conv}}\left(p2\right)\oplus p\right]$ ;
15          $\hat{p}\leftarrow lw\odot p2$ ;
          /* $\mathrm{Seg}\left(\cdot,\cdot\right)$ End */
          /* $\mathrm{RT}\left(\dots,\cdot,\dots\right)$ Start */
16          $\hat{\vartriangle}_{1}\left[:,:,s\left(0\right):e\left(0\right),s\left(1\right):e\left(1\right)\right]\leftarrow\hat{p}$;
          /* $\mathrm{RT}\left(\dots,\cdot,\dots\right)$ End */
17
18         end for
19       $\hat{\vartriangle}_{2}\leftarrow\textnormal{{S-Mix}}\left(\hat{\vartriangle}_{1}\right)\oplus\hat{\vartriangle}_{1}$;
20       $\hat{\vartriangle}\leftarrow\textnormal{{C-Mix}}\left(\hat{\vartriangle}_{2}\right)\oplus\hat{\vartriangle}_{2}$;
21
22      end for
23
24   end if
25
26 end if
return $\hat{\vartriangle}$
Algorithm 1 Compound Motion Segmentation
Appendix E Contribution Statement
This work represents a collaborative effort among all authors, each contributing expertise from different perspectives. The specific contributions are as follows:
  • Wenbo Huang (Southeast University, China; Institute of Science Tokyo, Japan): Firstly proposing the idea of applying RWKV in FSAR, implementing all code of Otter, designing wide-angle evaluation in § B, conducting all experiments, deriving all mathematical formulas, all data collection, all figure drawing, all table organizing, and completing original manuscript.
  • Jinghui Zhang (Southeast University, China): Providing experimental platform in China, supervision in China, writing polish mainly on logic of introduction, checking results, and funding acquisition.
  • Zhenghao Chen (The University of Newcastle, Australia): Writing polish mainly on authentic expression, clarifying the definition of wide-angle, amending mathematical formulas, checking experimental results, rebuttal assistance, and funding acquisition.
  • Guang Li (Hokkaido University, Japan): Refining idea of Otter, writing polish mainly on descriptions of results, rebuttal assistance, and funding acquisition.
  • Lei Zhang (Nanjing Normal University, China): Proposing the extra experiments on real wide-angle dataset VideoBadminton, guiding CAM visualization, rebuttal assistance, and funding acquisition.
  • Yang Cao (Institute of Science Tokyo, Japan): Verification of the overall structure, providing experimental platform in Japan, supervision in Japan, rebuttal assistance, and funding acquisition.
  • Fang Dong (Southeast University, China): Funding acquisition.
  • Takahiro Ogawa (Hokkaido University, Japan): Writing polish mainly on numerous details, guiding t-SNE visualization, rebuttal assistance, and funding acquisition.
  • Miki Haseyama (Hokkaido University, Japan): Funding acquisition.
Acknowledgments
The authors would like to appreciate all participants of peer review and cloud servers provided by Paratera Ltd. Wenbo Huang sincerely thanks those who offered companionship and encouragement during the most challenging times, even though life has since taken everyone on different paths. This work is supported by Science and Technology Major Special Program of Jiangsu Province under Grant No. BG2024028; Jiangsu Provincial Frontier Technology Research and Development Program under Grant No. BF2024070; National Natural Science Foundation of China under Grant Nos. 62472094, 62572119, 62232004, 62373194; Shenzhen Science and Technology Program under Grant No. KJZD20240903100814018; Jiangsu Provincial Key Laboratory of Network and Information Security under Grant No. BM2003201; Key Laboratory of Computer Network and Information Integration (Ministry of Education, China) under Grant No. 93K-9; Collaborative Innovation Center of Novel Software Technology and Industrialization; Collaborative Innovation Center of Wireless Communications Technology; the Fundamental Research Funds for the Central Universities; JSPS KAKENHI Nos. JP23K21676, JP24K02942, JP24K23849, JP25K21218, JP23K24851; JST PRESTO Grant No. JPMJPR23P5; JST CREST Grant No. JPMJCR21M2; JST NEXUS Grant No. JPMJNX25C4; and Startup Funds from The University of Newcastle, Australia.

### v3

@@
 Extra evaluation on the VideoBadminton dataset further validates the superiority of Otter in wide-angle FSAR.
-Codehttps://github.com/wenbohuang1002/Otter
 1 Introduction
@@
 Therefore, FSAR has diverse real-world applications, including health monitoring and motion analysis (yan2023feature; wang2023openoccupancy).
-However, recognizing similar actions under regular viewpoint is a non-trival problem in FSAR.
+However, recognizing similar actions under regular viewpoint is a non-trivial problem in FSAR.
 For instance, distinguishing “indoor climbing” and “construction working” is challenging, as subjects exhibit similar actions against a wall.
 To mitigate this issue, wide-angle videos provide contextual background, such as a “climbing wall” or a “construction site”, expressing actions within specific scenarios more accurately.
-According to established definitions (lai2021correcting; zhang2025madcow), wide-angle videos with a greater field of view (FoV) are widespread [^1] This work adopts the widely accepted definition of wide-angle as FoV exceeding 80..
+According to established definitions (lai2021correcting; zhang2025madcow), wide-angle videos with a greater field of view (FoV) are widespread***This work adopts the widely accepted definition of wide-angle as FoV exceeding 80..
 FoV estimation (lee2021ctrl; hold2023perceptual) on popular FSAR benchmarks further reveals that approximately 35% of samples per dataset fall into this category, yet remain unexplored.
@@
 $3\times 224\times 224$ random crops and horizontal flipping data augmentation is applied during training while only the center crop is utilized in testing.
-As an exception, horizontal flipping is absent in SSv2 because of many actions with horizontal direction such as “Pulling S from left to right222“S” means “something”.”.
+As an exception, horizontal flipping is absent in SSv2 because of many actions with horizontal direction such as “Pulling S from left to right“S” means “something”.”.
 Implementation Details and Evaluation Metrics
@@
 Wenbo Huang sincerely thanks those who offered companionship and encouragement during the most challenging times, even though life has since taken everyone on different paths.
-This work is supported by Science and Technology Major Special Program of Jiangsu Province under Grant No.
-BG2024028; Jiangsu Provincial Frontier Technology Research and Development Program under Grant No.
-BF2024070; National Natural Science Foundation of China under Grant Nos.
-62472094, 62572119, 62232004, 62373194; Shenzhen Science and Technology Program under Grant No.
-KJZD20240903100814018; Jiangsu Provincial Key Laboratory of Network and Information Security under Grant No.
+This work is supported by Frontier Technologies Research and Development Program of Jiangsu under Grant No.
+BF2024070; National Natural Science Foundation of China under Grants Nos.
+62472094, 62072099, 62232004, 62373194, 62276063; Jiangsu Provincial Key Laboratory of Network and Information Security under Grant No.
 BM2003201; Key Laboratory of Computer Network and Information Integration (Ministry of Education, China) under Grant No.
-93K-9; Collaborative Innovation Center of Novel Software Technology and Industrialization; Collaborative Innovation Center of Wireless Communications Technology; the Fundamental Research Funds for the Central Universities; JSPS KAKENHI Nos.
+93K-9; the Fundamental Research Funds for the Central Universities; JSPS KAKENHI Nos.
 JP23K21676, JP24K02942, JP24K23849, JP25K21218, JP23K24851; JST PRESTO Grant No.
@@
 JPMJNX25C4; and Startup Funds from The University of Newcastle, Australia.
+Codehttps://github.com/wenbohuang1002/Otter